40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 6032     Entry time: Tue Nov 29 02:09:15 2011
Author: Zach 
Type: Update 
Category: RF System 
Subject: EOM temp stabilization fixed 

I inspected the temperature stabilization circuit today to see why it wasn't working. It didn't make sense that it just kept railing the heater even though the error signal was negative (which should turn the heater current OFF).

It turns out that the LF356 (FET-input op amp) that serves as the filter stage for the heater driver was broken---I measured a full, railed positive output even though the input was negative. We didn't have any more LF356s, so I replaced it with an OPA604 (Burr-Brown FET-input), and all seemed to work fine.

Since we were having too much digitization noise, I also added a temperature monitor buffer using a non-inverting OP27 circuit with G=99. This makes the overall calibration ~7.6 V/K into the ADC.

Below is a time series showing that it is working. The circuit was turned on near the beginning, and you can see that the heater is railed at +10V until shortly after the error signal goes negative, at which point it adjusts and ultimately approaches a steady-state value of ~7.8V.

EOM_temp.png

I have no figures to demonstrate this, but it seems that even with this 100x increase in monitor gain, the error signal is still below the ADC noise level. This could be because the ambient temperature fluctuations are just that small on timescales of less than a few hours. I'm not sure if we really need to be able to see the temperature noise level above a few mHz, but if we do we will have to find some way to increase our dynamic readout range. 

Also, Koji told me where he stashed the nice protoboards, so I will get onto transferring this circuit onto one ASAP. Since it is working now, I think I'll leave it until after the TAC.

ELOG V3.1.3-